OpenStudy (anonymous):

5x-2y=-24 -4y=-48-10x if the system cannot be solved state that it cannot be solved and explain why

OpenStudy (turingtest):

let's do a little rearranging on that second equation: 5x-2y=-24 10x-4y=-48 now divide the second equation by 2 what are the two equations now?

OpenStudy (anonymous):

the second row or the whole thing divide by 2?

OpenStudy (turingtest):

second "row" I suppose not the first equation, only the second but then write then both out again after so you can compare them

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

would it be 5x-2=-24 and the other 1 is 5x-2=-24

OpenStudy (turingtest):

you dropped the y write them one over the other; it's more clear that way

OpenStudy (turingtest):

5x-2y=-24 5x-2y=-24 hm... they are both the same what does that mean for the solutions?

OpenStudy (anonymous):

that it doesnt have 1?

OpenStudy (turingtest):

if there is no solution, this system should lead to a contradiction let's see if that's true...

OpenStudy (anonymous):

but i have a question y did u divide by 2? like how did u know?

OpenStudy (turingtest):

I noticed that the first equation is a multiple of the second you should always look for that kind of thing in systems ask yourself "what can I multiply these equations by in order to get rid of (at least) one of the variables" it does take a little practice to recognize these things of course

OpenStudy (turingtest):

so back to where we were: 5x-2y=-24 5x-2y=-24 do the same as before and subtract one equation from the other make sure to write out both sides what do you get by subtracting one eqn from the other

OpenStudy (anonymous):

like just subtract dont multiply

OpenStudy (turingtest):

just subtract the second equation from the first keep the left and right sides of the equations separate, as we did in the last problem

OpenStudy (anonymous):

if i just subtract both of them i get 5x-2y=-72 am i right?

OpenStudy (turingtest):

not quite, be careful with those negative signs and don't forget to subtract the left side as well; that is essential 5x-2y=-24 5x-2y=-24 ----------- 5x-2y-(5x-2y)=-24-(-24) ^^^simplify both sides

OpenStudy (anonymous):

ok so it would be 0? right

OpenStudy (turingtest):

careful: write out both sides or else that tells you nothing about the solutions

OpenStudy (turingtest):

that = sign can't just disappear like that

OpenStudy (anonymous):

so wat would it be?

OpenStudy (turingtest):

what is on the left ? 5x-2y-(5x-2y)=?

OpenStudy (anonymous):

10x and 4y?

OpenStudy (turingtest):

5x-2y-(5x-2y)=5x-2y-5x+2y=0 what about the other side?

OpenStudy (anonymous):

it would be 0 too

OpenStudy (turingtest):

so then we can write 0=0 that's what I wanted from you, because now I get to ask you: how often is that statement true?

OpenStudy (anonymous):

o lol the statement but it doesnt have a solution right?

OpenStudy (turingtest):

0=0 is \(always\) true that means that there are \(infinite\) solutions whenever the system can be reduced to a contradiction, like 1=2 then there are no solutions whenever we can reduce the system to something that is always true, like 1=1 then there are infinite soutions

OpenStudy (turingtest):

solutions*

OpenStudy (anonymous):

ooo ok i get it thank you soo much ur really good in explaining :)

OpenStudy (turingtest):

thanks I try and thanks for listening :)

OpenStudy (anonymous):

well ur doing good ;) thnx